OpenStudy (anonymous):

Using the law of cosines, in triangle DEF, if e=18 yd, d= 10 yd, f= 22yd, find angle D.

myininaya (myininaya):

 We are given SSS this means we are given all 3 sides this means we indeed have to use the law of cosines as suggested. Pretend we have a triangle with the following labels for the angles of that triangle: ABC (in that order that they occur) The sides being a,b, and c. Where a is opp of A where b is opp of B where c is opp of C So the law of cosines for it would be a^2=b^2+c^2-2bccos(A) b^2=a^2+c^2-2accos(B) c^2=a^2+b^2-2abcos(C)